OpenStudy (anonymous):

what is the slope (3, -2)?

OpenStudy (anonymous):

one point does not have a slope a line has a slope

OpenStudy (anonymous):

The whole question is: Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2.

OpenStudy (anonymous):

you have to use the point slope formula.

OpenStudy (anonymous):

And these are the choices: y = 2x - 8 y = 2x - 4 y = 2x + 7 y = 2x + 5

OpenStudy (anonymous):

Put it into Point-slope form first. Then it'll be easier to convert to slope-intercept. Point-slope: y - y1 = m(x - x1) Slope-intercept: y = mx + b

OpenStudy (anonymous):

plug in what you have and solve for the slope.

OpenStudy (unklerhaukus):